OpenStudy (anonymous):

find the derivative of (2x+1)ln(x^2+5)

zepdrix (zepdrix):

Oh my goodness... I've been spending the past 5 minutes trying to solve the integral... It's just a derivative.. Oh I'm silly :3 ok ok ok let's see. Looks like we'll need to start with the product rule.

OpenStudy (anonymous):

is it (fg)'= f' g + f g' ?

zepdrix (zepdrix):

Ah sorry I got distracted for a few minutes there D: Yes very good c:

zepdrix (zepdrix):

\[\large \frac{d}{dx}(2x+1)\ln(x^2+5)=\] \[\large \color{royalblue}{\frac{d}{dx}(2x+1)}\ln(x^2+5)+(2x+1)\color{royalblue}{\frac{d}{dx}\ln(x^2+5)}\] Ok understand the setup? we need to take the derivative of the blue pieces.

OpenStudy (anonymous):

i understand that and got up until the derivative ln(x^2+5). i don't understand how to get that

zepdrix (zepdrix):

Do you know the derivative of \(\ln x\)?

OpenStudy (anonymous):

1/x?

zepdrix (zepdrix):

yes, good. We'll do the same thing here, but we'll have to be careful about what our `x` is. And we'll also need to apply the chain rule. Here's a quick silly example just in case the concept is a little confusing.

zepdrix (zepdrix):

Whatever the `argument` is (the thing inside of the log), that entire contents gets put into the bottom. Then we apply the chain rule, multiplying by the derivative of the inner function.

zepdrix (zepdrix):

BAHHH my pictured got messed up, there should be a prime on that far right potato.

zepdrix (zepdrix):